### v1

# INDIC QA BENCHMARK: A Multilingual Benchmark to Evaluate Question Answering capability of LLMs for Indic Languages

Abstract
Large Language Models (LLMs) have demonstrated remarkable zero-shot and few-shot capabilities in unseen tasks, including context-grounded question answering (QA) in English. However, the evaluation of LLMs’ capabilities in non-English languages for context-based QA is limited by the scarcity of benchmarks in non-English languages. To address this gap, we introduce Indic-QA, the largest publicly available context-grounded question-answering dataset for 11 major Indian languages from two language families. The dataset comprises both extractive and abstractive question-answering tasks and includes existing datasets as well as English QA datasets translated into Indian languages. Additionally, we generate a synthetic dataset using the Gemini model to create question-answer pairs given a passage, which is then manually verified for quality assurance. We evaluate various multilingual Large Language Models and their instruction-fine-tuned variants on the benchmark and observe that their performance is subpar, particularly for low-resource languages. We hope that the release of this dataset will stimulate further research on the question-answering abilities of LLMs for low-resource languages.
INDIC QA BENCHMARK: A Multilingual Benchmark to Evaluate Question Answering capability of LLMs for Indic Languages
Abhishek kumar singh, Rudra Murthy§, Vishwajeet Kumar§ Jaydeep Sen§,Ganesh Ramakrishnan Indian Institute of Technology Bombay, India §IBM Research, India {abhisheksingh,ganesh}@cse.iitb.ac.in, {rmurthyv, vishk024, jaydesen}@in.ibm.com
"In a world deluged by irrelevant information, clarity is power."
– Yuval Noah Harari
1 Introduction
India has the largest global population, with almost 1.43 billion people. However, many of the major languages in India are considered low-resource by the natural language processing (NLP) community. These languages are not as well-represented in NLP compared to English because of the lack of high-quality datasets for pre-training, fine-tuning, and task-specific evaluations. In the field of NLP, Large Language Models (LLMs) have been pre-trained on vast amounts of textual data. Despite their extensive training, these models frequently yield inaccurate results in tasks such as question answering, largely due to their limited contextual comprehension and uncertainties surrounding their parametric knowledge. Researchers have addressed this issue by utilizing systems like Retrieval-Augmented Generation (RAG). RAG retrieves relevant text from a large corpus, whether it’s from pre-training data or real-world data, and uses it as the context for the query. While retrieval is important, the generator component, which finds the exact and correct answer without producing inaccurate outputs, is equally crucial, especially when the answer is not present in the retrieved text.Gao et al. (2023).
Figure 1: One instance from the evaluation benchmark in Hindi languages.
In a Retrieval-Augmented Generation (RAG) system, there are two distinct components: the retriever and the generator. Each component is evaluated separately. The retriever’s performance is assessed through tasks such as passage reranking, utilizing datasets like Thakur et al. (2021), where the retrieval model is trained to rank paragraphs based on their relevance to a given query. On the other hand, the generator’s evaluation focuses on context-grounded question answering. In RAG setups, context is obtained through retrieval, necessitating that the dataset structure includes triples comprising context, question, and answer. The model’s objective is to generate an answer to the query based on the provided context.
What distinguishes our benchmark from other existing Multilingual Indic context-grounded question-answering Benchmarks? There are numerous context-grounded question-answering benchmarks available for high-resource languages like English. However, there are very few benchmarks available for Indic languages, and those that do exist often lack domain diversity and are limited in size. To address these gaps, we developed the INDIC QA BENCHMARK, which not only includes a large number of data instances but also spans a wide range of domains, including geography, Indian culture, news, and more. Given the existing domain diversity in English datasets Lee et al. (2019); Bonifacio et al. (2021); Rajpurkar et al. (2016), we decided to translate these datasets into Indic languages. While there are many extractive question-answering datasets, there is a scarcity of abstractive question-answering datasets where the answer may not be explicitly present in the text. To address this, we sampled numerous Wikipedia and Common Crawl pages, focusing on paragraphs rich in cultural nuances and domain diversity. Using these paragraphs, we employed large language models (LLMs) to generate QA pairs, thus creating a comprehensive and culturally diverse benchmark for Indic languages.
In our observations, we have noticed that directly using a base pre-trained model does not initially predict or output answers effectively. Even when it does find an answer, the output is often incorrect and illogical, alongside the ground truth. However, after few-shot prompting, the model produces better-formatted answers and causes the model to start searching for exact answers in single words or phrases from the paragraph, for generation. We summarize the key contributions of this paper as follows:
  • INDIC QA BENCHMARK: a multilingual evaluation benchmark for evaluating Indic Question-Answering/Generative capability of large language models, all these languages are low-resource language and lack a proper benchmark for multi-domain generative tasks.
  • We critically evaluate some of the most esteemed Language Model Models (LLMs) for Indic languages of equivalent size, meticulously comparing their performance on our proposed benchmark to determine their Question-Answering (QA) skills.
Datasets As Bn Gu Hi Kn Ml Mr Od Pa Ta Te
Hindi Squad 3099 3107 3371 4734 3068 2926 3165 3079 3469 2743 2955
NQ Open 1462 1483 1570 1842 1447 1420 1511 1451 1570 1331 1420
Chaii 339 351 394 746 351 328 373 305 388 325 361
Indic QA 1789 1763 1369 1547 1517 1589 1604 1680 1542 1804 1734
XSquad 1190 1190 1190 1190 1190 1190 1190 1190 1190 1190 1190
XORQA 537 538 532 537 534 533 529 529 531 537 538
MLQA 2362 2403 2718 4918 2299 2128 2433 2370 2730 2129 2291
Synthetic MCQA 296 284 358 593 288 243 330 296 385 243 266
MS Marco 29724 30089 31741 35735 29212 28528 30180 30073 32032 27197 28995
Llama Index 1158 1312 1333 1384 1310 1250 1316 1263 1175 1258 1306
Table 1: Indic-QA Dataset Statistics. Indic-QA benchmark is a compilation of existing datasets, English datasets translated to Indian languages, and, synthetic dataset generated using Gemini. The dataset comprises Extractive Question Answering and Generative Question Answering ({*}). As: Assamese, Bn: Bengali, Gu: Gujarati, Hi: Hindi, Kn: Kannada, Ml:Malayalam, Mr: Marathi, Od: Odia, Pa: Punjabi, Ta: Tamil, Te: Telugu. We have translated NQ Open, XORQA, Llama Index, and MS Marco datasets to Hindi. We have translated all the above datasets to the remaining ten Indian languages (underline data instances were already present in the referred language.)
2 Related Work
In the realm of context-grounded question answering (QA), significant research has been conducted in both English and Indian languages. This task involves presenting a question along with a contextual paragraph to the model, which then extracts the phrase from the paragraph. Various benchmarks Dzendzik et al. (2021); Rajpurkar et al. (2016, 2018) have been established for this task, with encoder-only transformer models proving effective in extracting the span containing the answer from the paragraph. The Indic QA community has demonstrated remarkable performance using models like XLM-RoBERTaConneau et al. (2019) and others, particularly for multilingual Indian languages. They have a rich dataset to showcase their benchmarks, including SQuAD Rajpurkar et al. (2016) for English, along with its translated version in Hindi. Additionally, instead of translation, there are datasets specifically designed for evaluating benchmarks in Hindi, such as the Chaii  [^1] https://www.kaggle.com/competitions/chaii-hindi-and-tamil-question-answering. dataset and IndicQA, which are also discussed in this survey paper Kolhatkar and Verma (2023). Although there are a few benchmarks for Indic Question Answering, they lack extensive domain coverage, which is crucial for evaluating the robustness of models. In contrast, English benchmarks encompass a wide range of domain-specific datasets such as Resources like the llama Index Liu (2022) highlight that selecting the appropriate evaluation dataset is challenging and highly dependent on the specific use case. Academic benchmarks such as BEIR and HotpotQA often fail to generalize across different use cases. For example, parameters that work well on certain data domains (e.g., SEC filings) may not perform as effectively on others (e.g., research papers). This challenge led them to the create a dataset hub specifically designed for evaluating RAG systems, encompassing a wide range of domains including research papers, blockchain articles, and code articles.
Additionally, NQ Open Lee et al. (2019) contains a wealth of Wikipedia content across various domains, and MS MARCO Bonifacio et al. (2021) features questions sampled from real-world user searches with contexts derived from web documents. The diversity of user queries leads to a broad range of content, making MS MARCO highly versatile. Although initially intended for a different task, we adopted this dataset for our purposes. Hence to address the lack of existing Indic QA benchmark datasets, we translated and adapted several commonly used English QA datasets into 11 Indic languages. This approach provides a more comprehensive and robust evaluation framework for Indic Question Answering models. By leveraging these datasets, we aim to offer a diverse and extensive evaluation resource, enhancing the development and assessment of QA models in Indic languages.
Figure 2: Workflow of synthetic data creation, in the fig. LLM used is Gemini-pro model.
3 Benchmarks
The primary focus of this work is on context-based QA, where the answer is entirely or partially found within the given context. The datasets utilized in this study were tailored to facilitate this task, with each instance composed of triples consisting of a context, a question, and an answer. This section provides a detailed description of the methodology used to create or modify the existing dataset for our task.
3.1 Datasets
In this section, we provide a catalog of the datasets constituting this benchmark, complete with a thorough exposition of their original accessibility and the modifications we have implemented. These datasets are either pre-existing or have been released as part of this work. Following is a detailed description of each dataset.
  1. 1.
    Hindi SQuAD: This dataset is a translated version of the original SQuAD Rajpurkar et al. (2016) into Hindi. It consists of nearly 5,000 instances, translated using the Google Translate API. We translated that from Hindi to other Indic languages.
  2. 2.
    XQuAD: XQuAD (Cross-lingual Question Answering Dataset) Artetxe et al. (2019) serves as a benchmark for assessing the performance of cross-lingual question answering. It comprises 240 paragraphs and 1190 question-answer pairs, extracted from the development set of SQuAD v1.1 Rajpurkar et al. (2016). The dataset includes professional translations of these pairs into ten languages, But we have Used Singh et al. (2024) version of XQuAD because they manually translated to all Indic languages.
  3. 3.
    ChaII Dataset Thirumala and Ferracane (2022): This question-answering dataset features context-question-answer triples in Hindi and Tamil, gathered directly without translation. Created by expert data annotators who are native speakers, the dataset presents a realistic information-seeking task focused on predicting answers to genuine questions about Wikipedia articles. It was used in a Kaggle challenge and includes 1104 questions in Hindi and Tamil, we used the Hindi part of the data and translated it to 10 other Indian languages.
  4. 4.
    Indic QA Doddapaneni et al. (2022): This dataset is a manually curated cloze-style reading comprehension dataset designed for evaluating question-answering models in 10 Indic languages Since this dataset doesn’t have Gujarati translation we translated it from Hindi to Gujarati and validated the translation as described in the [3.2] section.
  5. 5.
    MLQA Lewis et al. (2019): MLQA (MultiLingual Question Answering) is a benchmark dataset for evaluating cross-lingual question answering performance. We have used the MLQA test set for benchmarking purposes, the test set contains 4918 triples of the form (context, question, answer) all available in Hindi, hence we translated this triplet from Hindi to 10 other Indian languages.
  6. 6.
    MS Marco Bonifacio et al. (2021): Microsoft Machine Reading Comprehension (MS MARCO) is a collection of large-scale datasets designed for deep learning applications related to search. The questions in MS Marco are sampled from real, anonymized user queries. The context passages, from which the answers are derived, are extracted from real web documents using the latest version of the Bing search engine. We initially considered adapting the multilingual version of the MS MARCO passage ranking dataset (mMarco) for our setting. However, since mMarco lacks a test set, we opted to use the MS MARCO test set, which contains 100k instances, each including a query and a set of passages, among which only one is relevant to the query. We filtered out instances without any relevant passages, resulting in a dataset of 55k instances.
    We then translated this dataset from English to Hindi. After applying certain filtering conditions, we translated the Hindi dataset into 10 other Indian languages. The exact steps are detailed in [3.2]. The final dataset now includes the question, the source document, and the corresponding answer, and is available in 11 Indian languages.
  7. 7.
    NQ-Open trans Lee et al. (2019): The NQ-Open task is an open-domain question-answering benchmark derived from Natural Questions. The objective is to predict an English answer string for a given English question, with all questions answerable using the contents of English Wikipedia. Initially, the dataset was entirely in English, with context, question, and answer all in English. The context often included tables scraped from HTML pages of Wikipedia, resulting in numerous HTML tags. To clean the dataset, we removed all triples where the context contained a table and eliminated all other HTML tags from the remaining examples. In this modified dataset, the fields include the source document (the entire Wikipedia page), the long answer (a paragraph from the page containing the answer), and the exact phrase or word from that paragraph as the short answer. We modified the long answer to serve as the context and the short answer as the answer for the corresponding question. and Since after all this modification dataset was in English we translated that to other Indian languages.
  8. 8.
    XORQA Asai et al. (2020): Cross-lingual Open Retrieval Question Answering (XOR QA) consists of three tasks involving cross-lingual document retrieval from both multilingual and English resources. This dataset was subsequently translated into other Indian languages by Singh et al. (2024). We utilized the same since it was cross-lingual data, the context was in English while the questions and answers were in other languages. To adapt it to our setting, we translated the context into various Indian languages.
  9. 9.
    LLama Index Liu (2022): The dataset includes question-answer pairs along with source context, serving as an evaluation tool for the RAG pipeline. We observed that some contexts were insufficient to answer the questions effectively. To address this, we applied the BGE-M3 Chen et al. (2024) algorithm to measure the similarity between the context and the query, using a threshold of 0.43 to determine if a question could be answered adequately based on the context. Post filtering we translated the resulting context, question, and answer triplets into Hindi and Hindi other Indian languages.
  10. 10.
    Synthetic Data: This dataset is introduced as part of this study. We employed the Gemini model Team et al. (2023) to generate question-answer pairs based on provided contexts. To achieve this, we sampled a diverse set of Hindi contexts from sources such as Wikipedia, storybooks, Indian news articles, and paragraphs from competitive exams. We then prompted the model with these context paragraphs to generate abstractive question-answer pairs, framing the task as a generative QA task. Subsequently, this dataset was translated into other languages and verified by language experts, the whole workflow process can be found 2.
3.2 Data Curation Methodology
In light of the approaches discussed previously in Section 2, context-grounded question-answering datasets can generally be categorized into two types: abstractive and extractive. While there are many extractive datasets available in high-resource languages, a few extractive datasets available for Indian languages lack diversity in domains and question types, limiting their utility for benchmarking. Hence, we extended the benchmark suite available in English to these Indian languages by translating. We utilized Indic Transv2  [^2] https://github.com/AI4Bharat/IndicTrans2 Gala et al. (2023) for translation, an open-source transformer-based multilingual NMT model that supports high-quality translations across all the 22 scheduled Indian languages. We segmented the context paragraph into sentences using the Spacy library, translated each sentence, and then recombined them. This approach yielded better translation results, and importantly, the model did not lose context when translating, thus preserving the coherence of the text.
In the list of datasets for benchmarking, some are available only in English (e.g., NQ-open, ORQA, llama index,MS-Marco), while others are available in both English and Hindi (e.g., Hindi SQuAD, CHAII, MLQA, Synthetic data). Additionally, a few datasets (e.g., IndicQA, XSQuAD) are also available in all 10 or 11 languages with verified translations. For all the datasets not found in the respective language, we translated them and applied the filtering methods discussed below.
To assess the quality of our translations, we first translated each dataset from the source language to the target language, then back-translated it from the target language to the source language. We calculated the CHRF and CHRF++ scores between the original and back-translated sentences, applying a threshold on these metrics to filter the instances. Additionally, we manually verified a subset of the filtered data to ensure accuracy. For the translation process, we initially translated the English data directly into Hindi. After filtering the data, we then translated it from Hindi to other Indian languages, rather than directly from English. This approach was based on our observation that the translation quality from Hindi to other Indian languages was superior. The improved quality can be attributed to the linguistic similarities within the same language family, including morphology, syntax, and grammar.
Languages Bloom Gemma Llama-3 Openhathi
Ext Gen Ext Gen Ext Gen Ext Gen
As 13.86 5.04 21.70 4.92 19.23 3.85 1.14 0.02
Bn 17.84 7.88 21.15 4.93 19.14 4.22 1.49 0.03
Gu 13.27 7.15 24.90 6.36 20.27 5.35 4.07 0.05
Hi 21.69 7.65 34.37 5.83 41.53 8.08 11.83 0.004
Kn 15.63 7.40 24.47 5.63 20.39 4.52 0.31 0.00
Ml 19.22 8.36 25.20 5.20 22.80 5.83 0.68 0.002
Mr 15.12 6.69 23.96 5.43 36.12 7.56 2.07 0.00
Od 11.11 5.88 9.06 2.58 11.23 5.31 0.26 0.00
Pa 15.60 6.87 28.43 6.28 21.96 5.92 1.31 0.00
Ta 19.96 9.07 22.45 5.30 19.74 4.99 0.70 0.00
Te 16.07 7.89 23.93 7.32 13.59 3.17 0.53 0.00
Average 16.31 7.26 23.60 5.43 22.36 5.35 2.22 0.0096
Table 2: Performance of various Large Language Models on Zero-Shot INDIC QA BENCHMARK. We report the average numbers across span-extraction datasets and abstractive question-answering datasets.
Languages Bloomz Gemma-Instruct Llama-3-Instruct Airavata Aya-8-Instruct
Ext Gen Ext Gen Ext Gen Ext Gen Ext Gen
As 38.69 7.36 11.32 9.97 15.22 3.42 3.70 3.41 19.39 7.19
Bn 44.75 8.27 13.03 11.66 17.89 3.90 6.58 4.38 29.79 8.27
Gu 49.06 9.24 7.61 7.74 12.50 2.90 5.15 2.87 26.48 5.82
Hi 62.88 9.06 18.54 7.61 14.95 7.21 44.35 7.21 55.17 8.02
Kn 40.10 8.27 11.94 8.80 15.88 3.13 1.39 1.07 16.22 8.62
Ml 43.84 8.55 9.33 5.28 13.67 2.32 7.55 6.80 30.35 8.55
Mr 50.03 8.09 14.15 10.95 13.84 1.87 14.15 5.91 35.83 9.78
Od 37.10 8.78 1.61 0.72 5.59 1.74 1.78 1.27 15.77 5.79
Pa 53.37 9.11 10.78 8.23 18.07 4.72 3.53 2.87 11.81 9.79
Ta 46.43 10.33 16.31 10.42 17.98 3.90 7.66 5.01 36.46 10.26
Te 44.42 8.98 10.85 7.53 15.04 4.13 4.43 3.48 18.82 7.61
Average 46.42 8.73 11.41 8.08 14.60 3.57 9.12 4.03 26.92 8.15
Table 3: Indic-RAG Dataset Statistics. Indic-RAG benchmark is a compilation of existing datasets, English datasets translated to Indian languages and Synthetic dataset generated using Gemini
Languages Bloom Gemma Llama-3 Openhathi
1 shot 3 shot 1 shot 3 shot 1 shot 3 shot 1 shot 3 shot
As 21.22 22.44 37.91 40.43 28.61 29.42 3.15 3.14
Bn 19.82 26.90 42.64 37.85 32.71 24.95 6.86 6.56
Gu 20.17 22.16 44.26 47.40 26.76 22.46 3.65 4.61
Hi 34.08 36.96 54.95 56.95 54.79 57.11 10.27 22.1
Kn 19.99 19.93 37.74 35.80 22.76 15.60 0.33 2.26
Ml 25.82 23.84 42.08 31.10 28.94 22.60 6.59 5.50
Mr 26.35 25.99 47.08 37.39 47.17 40.7 5.82 8.63
Od 15.72 15.68 16.66 22.05 13.37 5.39 0.37 3.32
Pa 23.18 23.98 42.18 47.17 31.03 30.62 2.00 3.52
Ta 26.80 26.97 43.27 43.90 28.66 27.06 4.51 4.94
Te 20.83 25.22 39.90 43.62 25.37 20.86 3.09 2.76
Table 4: Performance of various Large Language Models on few-Shot INDIC QA BENCHMARK. We report the average numbers across span-extraction datasets and question-answering datasets.
4 Experiment Setup
We conducted a series of experiments to evaluate existing LLM’s performance, utilizing the NVIDIA RTX A100 both 40Gb and 80Gb variants for our computational needs. Our computational needs signify GPUs both for Translation and evaluation over the models. For inference or evaluation we utilized VLLM Kwon et al. (2023) which is an open-source library that supports LLM inference efficiently.
We evaluate the following LLMs on our benchmark: Open Hathi and its Instruction Finetuned variant (IFV) known as Airavata Gala et al. (2024), Bloom Le Scao et al. (2022) and its IFV named Bloomz, GemmaTeam et al. (2024), and its instruction fine-tuned variant Gemma-IT. Open Hathi [^3] https://www.sarvam.ai/blog/announcing-openhathi-series (7B parameter model), which was created through continual pre-training on the LLaMA-2 model Touvron et al. (2023). Airavata Gala et al. (2024) (7B parameter model) is an instruction fine-tuned version of OpenHathi. Both OpenHathi and Airavata are specifically trained for Hindi.
Gemma and Gemma-IT (7B parameter models) [^4] https://ai.google.dev/gemma/docs, which were released by Google. Though these models are not specifically trained for Indian languages, they exhibit multilingual capability. Aya-8B Aryabumi et al. (2024) is another instruction-tuned model designed explicitly for multilingual languages. LLaMA 3 and LLaMA 3 Instruct models  [^5] https://ai.meta.com/blog/meta-llama-3/(8B parameter model), part of the LLaMA family. Llama-3 has seen data from around 30 languages excluding English. This diverse set of models allows for a comprehensive evaluation of the strength and performance of our benchmarks across different architectures and training methods.
4.1 Evaluation Metrics
We borrow the key metrics of both Extractive and generative Question Answering as mentioned below:
1. F1 (macro-averaged) score The F1 score is a metric that represents the harmonic mean of precision and recall. The F1 score calculates the average similarity between predicted and actual answers by comparing the sets of words or tokens in the predicted and ground truth sentences.
2. Exact Match The Exact Match metric computes the percentage of instances that exactly match the ground truth. This metric is more stringent but evaluates the thoroughness of the model, as it results in either a true or false outcome for individual instances.
3. ROUGE(L) A Recall-Oriented Understudy for Gisting Evaluation metric is mostly used for evaluating summarization, we have used it to evaluate generative QA tasks.
5 Results and Analysis
Comparing base model performance and effect of Few shot: Table 2 shows the base LLMs’ performance in the zero-shot setting. The Gemma model excels in extractive question-answering tasks, surpassing the Bloom and Llama-3 base models. However, Llama-3 outperforms Gemma in Hindi, Marathi, and Odia languages. Bloom surpasses all models in abstractive question-answering tasks and for all languages. Notably, base models generally perform poorly on abstractive question-answering datasets compared to extractive ones.
We also evaluate the effect of in-context examples as reported in Table 4. As expected, using few-shot (1-shot and 3-shot) almost always improves over the zero-shot base model. However, we can spot some language-specific patterns where Bloom and Openhathi behave differently than Gemma and Llama-3. For example, for some languages such as Bn, Ml, Mr Gemma and Llama-3 show a significant drop with the increase in few shot examples, however, Bloom and Openhathi retain or even improve. We believe this is correlated with the availability of language-specific corpus and their utilization in training these models.
Effect of Instruction Finetuning: Table 3 shows the performance of instruction-finetuned models in our study. Instruction finetuning generally improves abstractive QA tasks across all models, but its impact on extractive QA varies. For instance, Gemma and Llama-3 perform better than Bloom and OpenHathi in their base models, but their instruction-finetuned variants do not show significant improvement. This is because these models were primarily instruction-finetuned on non-Indic languages, which compromises their generic multi-lingual ability during task-specific finetuning, leading to lower results.
On the other hand, OpenHathi was specifically trained on the Hindi language and so is its instruction finetuning variant Airavata. As a result, the performance of OpenHathi is significantly poor in all languages. Airavata benefits from further instruction fine-tuning on Hindi data and improves over OpenHathi for Hindi language but suffers poorly for other Indian languages. Bloomz produces the highest jump compared to Bloom and we hypothesize this is because a good portion of evaluation benchmark coming from generic-domain such wikipedia data has been seen by Bloomz during its training and instruction finetuning, making it a good choice for applications which aims to use common world knowledge.
Extractive vs Abstractive tasks: While it is clear that instruction finetuning helps more in abstractive QA tasks, both Table 2 and Table 3 shows positive correlation between the scores for extractive task and abstractive task across languages i.e. whenever numbers have improved for extractive QA tasks, it also improved for abstractive QA too. This is almost true for all the base and instruct variants of the models except Gemma, where Gemma instruct improves abstractive QA score but detoriates in extractive QA task. Careful analysis shows, generative task metrics change moderately between base models and their instruction finetuned variants. This is expected because generative metrics such as rogue(l) are more heuristic driven in nature and designed to ignore small variations, natural in non-deterministic text generation, unlike extractive metrics such as EM, F1. Thus generative metrics deviate in smaller scale than extractive metrics. However, the positive correlation between both the task metrics across models clearly establishes that the factors affecting the overall performance of the models shows similar signs for both extractive and generative tasks and hence improving one will likely improve the other as well.
On how to choose a model: Going by the results so far, one would pick BloomZ if the application needs only common world knowledge and needs a model which does well OOB. If there is use-case for which we have adequate Indic language finetuning data , it might be good to build over the world knowledge acquired by Gemma and Llama-3 and do instruction finetuning on Indic languages to make it better suitable for abstractive QA task. If we are very specific about a certain niche domain in only Hindi language, where common world knowledge is not a pre-requisite, Airavata can be a good candidate given its focus on Hindi based training and improvements in both extractive and abstractive task with instruction finetuning.
6 Conclusion
In this paper, we release a benchmark for evaluating the grounded Question-Answering capabilities of existing Large Language Models (LLM). The benchmark comprises datasets testing both extractive and abstractive answering abilities. Our observations reveal that question-answering capability in LLMs can be enhanced through instruction-tuning of these models with target language data. We hope that this benchmark will aid researchers in improving LLMs’ grounded question-answering abilities.
7 Limitations
While our research aims to provide a challenging and comprehensive benchmark for evaluating the LLMs on Hindi QA task, there might be some limitations. 1. Availability of high quality datasets for Hindi is limited. Despite our best efforts to curate the benchmark from various sources, it might still have an inherent bias introduced during the data collection/translation process. 2. Although we conducted quality checks, there might be subjective interpretability issues with the translated datasets. 3. While we attempted to diversify across various domains, the benchmark may not depict the true performance in a completely unseen domain.
Appendix A Appendix
A.1 Miscellaneous
Figure 3: CHRF and CHRF++ scores computed after translating the NQ open dataset from English to Hindi and then back-translating from Hindi to English. The scores are calculated between the original and back-translated parts, with a threshold of 50 applied to the CHRF scores to filter the data.
Figure 4: Evaluation prompt used for the base model.
Figure 5: Evaluation prompt used for the instruction-tuned model. Most models utilize a chat format. This is an example of the prompt used for the Airavata model.
Figure 6: Base Model(Bloom) prediction on one of the instance of extractive QA dataset, it is evident from the fig. the base model sometimes output garbage along with correct answer.
Figure 7: Instruction tunned Model(Bloomz) prediction on one of the instance of extractive QA dataset.
Figure 8: Base Model(Bloom) prediction on one of the instance of Generative QA dataset
Figure 9: Instruction tunned Model(Bloomz) prediction on one of the instance of Generative QA dataset.
Figure 10: